# MedMax: Mixed-Modal Instruction Tuning for Training Biomedical Assistants

### v1

@@ 5 Experimental Setup
Post data creation, we intend to instruction-tune a mixed-modal foundation model on the MedMax data (§5.1). Subsequently, we will present the evaluation methods to assess the performance of our model across diverse skills (§5.2).
@@ 5.1 MedMax Mixed-Modal Model
In this work, we will instruction-tune an instantiation of the Chameleon-7B [54] mixed-modal foundation model. Specifically, it can understand and generate multimodal content including images and text. Primarily, it represents the raw images as discrete visual tokens using a VQGAN [17], and the text data into discrete text tokens using BPE tokenizer [48]. Subsequently, each instance in the training dataset is represented as a sequence of discrete tokens, and the model is trained to predict the next token based on the preceding tokens in the sequence, following an autoregressive objective. The model consists of a vocabulary size of $65536$ where $8192$ are visual tokens apart from beginning of image and end of image tokens. In addition, the vocabulary includes a reserved token ‘<reserved08706>’ that separates the instruction (context) from the response (output) for instruction-tuning. Post-tokenization, the entire MedMax data consists $1.7$B tokens where $0.7$B and $1$B are visual and text tokens, respectively.
While the Chameleon-7B model weights are publicly available, they were safety-tuned and support mixed-modal inputs and text-only output to be used for research purposes. [^2] https://ai.meta.com/blog/meta-fair-research-new-releases/ To unlock the mixed-modal output capabilities, Anole [15] selectively finetunes the output embeddings of the image tokens using high-quality images from LAION [47]. This strategy does not interfere with the input mixed-modal understanding and text-only output abilities of the original Chameleon model. Hence, we choose Anole-7B as our base model for mixed-modal instruction tuning in this work. In particular, we utilize low rank adaptation (LoRA) [21] for parameter-efficient finetuning of the base model. Further, we finetune the base model for $3$ epochs on the MedMax dataset. We provide more finetuning details in Appendix G.
While various mixed-modal models, such as Transfusion [76] and Monoformer [75], offer different approaches, we select Chameleon (or Anole) for its simplicity in utilizing a single autoregressive loss, avoiding the complexity of balancing multiple objectives like autoregressive and diffusion. However, we note that MedMax can enable finetuning of any of these mixed-modal models too.
@@ 5.2 Evaluation
Post-training, it is critical to evaluate the capabilities of the instruction-tuned mixed-modal model across diverse skills. We provide the summary of the tasks and datasets in our comprehensive evaluation suite in Table 2. We will present the individual components here:
@@
This task involves answering complex questions about the biomedical images from different domains and anatomical regions of the human body. Specifically, we include the test set of VQA-RAD (radiology), SLAKE (semantic knowledge over radiology), PathVQA (pathology), and the entire QuiltVQA (histopathology) dataset. These datasets ask closed-ended (yes/no) and open-ended questions that require one word, phrase or sentence answer. Here, we use exact match to assess the accuracy of the models on the closed-ended questions. However, the evaluation on the open-ended questions is intrinsically harder due to the subjectivity of the answers. Prior works [33] utilize the ratio of the ground-truth tokens appear in the generated answer, but this is quite unreliable when there are no constraints on the response format [37]. For instance, if the ground-truth is ‘bronchi’ and predicted answer is ‘lung’, then the metric will penalize the predicted answer. To avoid such challenges, we utilize an LLM (GPT-4o-mini) that compares the predicted answer against the ground-truth answer to decide where the model outputs are reliable or not. For each open-ended question, it gives a score of 0 or 1. We provide the evaluation template in Appendix Table 12.
Additionally, we also include medical VQA with multiple-choice questions datasets such as test set of the PMC-VQA (diverse biomedical domains), validation set of PathMMU (pathology) [52], and ProbMed (radiology) [65] dataset. In addition, we assess the performance of an hidden split of $1000$ questions from the OmniMedVQA [22] dataset. Overall, we perform evaluations on 12 VQA tasks across diverse biomedical domains, skills, and question formats.
@@
Here, we aim to compare the capability of the MedMax model and the base Chameleon (Anole) modle to caption as well as generate biomedical images from different domains. In total, we collect $1200$ instances from the testing split of PMC-OA (400), MIMIC-CXR (400), and unseen split of Quilt-1M (400) datasets. In particular, half of the dataset will be used for captioning evaluation and the other half will be used for generation evaluation. Similar to [13], we extract the impressions (summary of the findings) from the report data and treat them as the ground-truth captions for the associated chest radiographs in the MIMIC-CXR dataset. In principle, this will highlight the medical report understanding and report-conditioned image generation capability of our model. Subsequently, we will compute the BioMedCLIPScore [72] to assess the closeness between the input image and predicted caption (and vice-versa). We present the details for model inference in Appendix E.2.
Biomedical Visual Chatbot: LLaVA-Med introduced a visual chatbot evaluation where the task of the model is to answer $193$ novel questions about $50$ unseen biomedical images. Specifically, the questions belong to two category: conversation and detailed description of the images. Subsequently, an LLM scores the predicted answer and the GPT-4 written reference answer out of 10 conditioned on the question, image caption and additional image context. Finally, we compute the average relative prediction score as the ratio of the score for predicted answer and score for the reference answer. In our work, we leverage GPT-4o-mini for the scoring the generated responses due to its cost-effectiveness and better capabilities than GPT-4.333https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/
Biomedical Multimodal Generation: With the development of mixed-modal biomedical assistants, it will be critical to assess their performance quantitatively. To this, we reserve $500$ instances of the MedMax-Instruct data for testing and are never exposed during the training process. We note that each instance of this dataset consists of text input and multimodal (text and image) output. For a given text query, we prompt the base (or finetuned) model to generate interleaved response. Subsequently, we compare the text content in the predicted multimodal response with the reference text response using LLM-scoring, identical to the biomedical visual chat evaluation. In addition, we compare the generated image with the reference image using the image-image similarity score from the BioMedCLIP model. We present the evaluation templates and inference details in Appendix E.2.
We performed a data contamination analysis to verify the independence between our MedMax dataset and the test sets used in all downstream tasks. Our examination revealed no exact matches between image-text pairs across these datasets, confirming the absence of data leakage.
@@ 6 Experiments
In this section, we will present the downstream performance of the instruction-tuned mixed-modal model with the MedMax data. In particular, we provide the results for VQA tasks, image captioning and generation, multimodal generation, and visual chat (§6.1). To study the dataset scaling, we will analyze the performance by finetuning the mixed-modal on different subsets of MedMax6.2). Finally, we conduct ablation studies to understand the impact of the specialized visual encoder, and role of specific data subsets on downstream performance (§6.3).
Table 3: Performance of the MedMax model and baselines on the downstream VQA tasks. We find that the MedMax mixed-modal model outperforms closed as well as open multimodal models on twelve VQA datasets. This highlights that the model can generalize well to unseen instances and tasks ranging across several biomedical domains.
@@ 6.1 Main results
Here, we will present our empirical findings on the diverse downstream tasks.
6.1.1 Biomedical Visual Question Answering
We compare the performance of the finetuned MedMax model against several vision-language models on the battery of the VQA datasets in our evaluation suite. Specifically, the baselines includes the open models: the mixed-modal Chameleon-7B [54], and multimodal input but text-only biomedical specific models: LLaVA Med-v1.5 [39], and HuatuoGPT-Vision-7B [71]. In addition, we consider two closed vision-language models: GPT-4o-mini and GPT-4o. We present our results in the Table 3.
@@ 6.1.1 Biomedical Visual Question Answering
Our experiments reveal that MedMax model outperforms the base model (Chameleon) by $26.1\%$ percentage points on the average VQA performance across twelve tasks. This highlights that MedMax is suitable for building biomedical specialized model from the base model via instruction-tuning. Moreover, we observe that it achieves the best average performance across diverse baselines. In particular, it beats the closed vision model, GPT-4o, by $18.3\%$ percentage points. It indicates that the MedMax model is the most capable multimodal model for biomedical VQA across open as well as closed models. Further, we highlight the task-specific performance on individual VQA tasks too.
@@
In addition, we observe that MedMax achieves the highest accuracy amongst the baselines for 7 out of 12 tasks. Interestingly, we find that Chameleon-7B achieves the highest QuiltVQA (Closed) but a low QuiltVQA (Open) accuracy which suggests that closed-ended questions are easier to solve for the base model. In addition, we observe that MedMax achieves a performance of $99.5\%$ on the unseen split of OmniMedVQA, highlighting that this dataset is not particularly challenging. The model demonstrates the ability to answer questions about diverse anatomical regions after being exposed to some in-distribution data from OmniMedVQA. Furthermore, we note that HuatuoGPT-Vision-7B outperforms other baselines on 3 out of the 12 datasets. This can be attributed to its high-quality fine-tuning dataset and the Qwen2 backbone [66], which was pretrained on 7 trillion text tokens. This backbone significantly enhances the model’s ability to comprehend diverse biomedical domains and formatting styles more effectively.
@@
Additionally, we observe that the MedMax model achieves performance improvements of $10.7\%$, $15\%$, and $23\%$ on the unseen QuiltVQA (Open), PathMMU, and ProbMed datasets, respectively. This underscores the MedMax model’s ability to generalize effectively to novel downstream tasks through exposure to expert knowledge, question formats, and biomedical domains via instruction-tuning. Furthermore, we find that the MedMax is competitive to the task-specific finetuning of the LLaVA med on VQA-RAD, SLAKE and PathVQA datasets (Appendix F). This indicates that the practitioners can utilize an unified MedMax model instead of hosting separate task-specific LLaVA med models. Overall, these results highlight at the high-quality of the MedMax dataset for finetuning the mixed-modal foundation models for biomedical assistants.
@@
6.1.2 Biomedical image captioning and generation
Here, we compare the capability of the MedMax model and base model to interpret and generate biomedical image data across different biomedical domains. We present the results in Figure 7. Our empirical findings that the MedMax consistently outperforms the base model on the biomedical image captioning as well as the image generation task. In particular, the MedMax outperforms Chameleon by achieving a relative gain of $28\%$, $33\%$ and $14\%$ on biomedical captioning for the images in the PMC-OA, Quilt, and MIMIC-CXR datasets, respectively. Additionally, the MedMax also outperforms the Chameleon by achieving a relative gain of $100\%$, $14\%$, and $50\%$ on the image generation tasks across PMC-OA, Quilt,and MIMIC-CXR captions, respectively.
Further, we compare the captioning abilities of our model against other models in Appendix H. In particular, we observe that the average BioMedCLIPScore for our model is better LLaVA Med-v1.5 and at par with baselines such as HuatuoGPT-Vision, GPT-4o-mini, and GPT-4o. However, we highlight that none of the those models can perform image generation natively, unlike MedMax and Chameleon models. Overall, our results demonstrate that the MedMax model excels at reasoning about images and generating realistic biomedical visuals aligned with user intent.
6.1.3 Biomedical Multimodal Generation
(a) Assessment of text content in MM output.
(b) Assessment of image content in MM output.
Figure 8: Performance on the multimodal generation task. Comparison between the performance of the MedMax and Chameleon mixed-modal model on the multimodal generation task. (a) We use LLM score to assess the quality of the generated data against the reference answer. (b) We use the image-image similarity score using BioMedCLIPScore to [72] compare the generated image and reference image. We find that MedMax finetuning improves the multimodal content generation capabilities for the biomedical domain.
Here, we aim to study the capability of the MedMax model to generate multimodal (interleaved image-text) content. We present the results in Figure 8. Our empirical findings highlight that the MedMax outperforms Chameleon by achieving a relative improvement of $25.2\%$ on the quality of the text content in the multimodal output, as measured by LLM score. Furthermore, we observe that the MedMax outperforms Chameleon by achieving a relative improvement of $31.5\%$ on the synthesized biomedical image in the multimodal output, as measured by the image-image similarity score using BioMedCLIP. This indicates that instruction tuning with MedMax data equips a mixed-modal model with strong multimodal generation capabilities in the biomedical domain. Overall, our results lay the foundation for future exploration in unlocking and evaluating multimodal generation capabilities with mixed-modal foundation models via instruction-tuning.
6.1.4 Biomedical Visual Chat
@@
Here, we study the ability of the MedMax model and other baselines including Chameleon, LLaVA-Med-v1, LLaVA-Med-v1.5, and HuatuoGPT-Vision to answer novel queries about the biomedical images in the PMC dataset. We present the results in Figure 9. We observe that the overall LLM score of MedMax model is higher than the base Chameleon model and LLaVA med-v1 by $34\%$ and $10.2\%$ percentage points. This indicates instruction-tuning enables strong visual chatting capability to the mixed-modal models. In addition, the MedMax model is quite competitive in comparison to the best-performing model, LLaVA Med-v1.5, with a difference of $2.1\%$ percentage points.
The fine-grained analysis indicates that the LLaVA med-v1.5 gains most of the performance on the conversation split of the dataset, while HuatuoGPT-Vision is better on the description split of the dataset. Such improvements in visual chat performance for LLaVA med-v1.5 and HuatuoGPT-Vision can be attributed to the quality of their language backbones, namely Mistral-v0.2-Instruct [25] and Qwen-2 [66], respectively. These backbones enhance the models’ ability to interpret human queries and generate high-quality textual outputs.We believe that future advancements in base mixed-modal models will further enhance performance on biomedical visual chatting using the MedMax dataset.
@@ 6.2 Data scaling
Figure 9: Performance on the visual chat task. Comparison between the visual chat performance of MedMax model and several baselines. We find that the chatting capabilities of our model is quite competitive, suggesting its ability to answer novel queries about biomedical images.
@@
6.3 Ablation Studies
@@ 6.3.1 Impact of Specific Data Subsets
6.3.2 Impact of Specialized Visual Encoder
@@ 8 Conclusion
In this work, we present MedMax, an instruction-tuning dataset designed for training mixed-modal foundation models. This dataset equips models with diverse multimodal biomedical capabilities across several domains. Notably, it is the first instruction-tuning dataset to enable multimodal generation capabilities for biomedical AI (MedMax-Instruct). Additionally, MedMax allows mixed-modal models to respond effectively to novel human instructions, functioning as capable biomedical assistants. Our experiments demonstrate that the MedMax-tuned model achieves competitive results across various downstream tasks, including biomedical visual question answering, multimodal generation, image captioning, image generation, and visual chatting. Overall, our work establishes a strong foundation for training the next generation of mixed-modal biomedical AI assistants.

### v2

@@ 4 Experimental Setup
Post data creation, we instruction-tune a mixed-modal foundation model on the MedMax4.1). Then, we present the evaluation framework for robust assessment (§4.2).
@@ 4.1 MedMax Mixed-Modal Model
In our work, we instruction-tune Anole-7B [9], an instantiation of the Chameleon-7B [34] mixed-modal foundation model that can natively understand and generate multimodal content. We use LoRA [14] for parameter-efficient finetuning of the model. Further, we finetune the base model for $3$ epochs on MedMax. We provide more details in Appendix G. [^2] While various mixed-modal models, such as Transfusion [69] and Monoformer [68], offer different approaches, we select Chameleon (or Anole) for its simplicity. Thus, we note that MedMax can enable finetuning of any of these mixed-modal models too.
@@ 4.2 Evaluation
While there are several multimodal biomedical datasets, there is a general lack of a comprehensive evaluation benchmark. To address this, we introduce a thorough evaluation suite to assess the capabilities of native multimodal models across various tasks and domains.
@@
We include the test set of VQA-RAD (radiology), SLAKE (semantic knowledge over radiology), PathVQA (pathology), and the entire QuiltVQA (histopathology) dataset. These datasets ask closed-ended (yes/no) and open-ended questions that require one word, phrase or sentence answer. Here, we use exact match to assess the accuracy of the models on the closed-ended questions. However, the evaluation on the open-ended questions is intrinsically harder due to the subjectivity of the answers. To avoid such challenges, we utilize an LLM (GPT-4o-mini) that compares the predicted answer against the ground-truth answer to decide where the model outputs are reliable or not. For each open-ended question, it gives a score of 0 or 1. We provide the evaluation template in Appendix Table 12. Additionally, we also include medical VQA with multiple-choice questions datasets such as test set of the PMC-VQA (diverse biomedical domains), validation set of PathMMU (pathology) [47], and ProbMed (radiology) [59] dataset. In addition, we assess the performance of an hidden split of $1000$ questions from the OmniMedVQA [15] dataset. Overall, we perform evaluations on twelve VQA tasks across diverse biomedical domains, skills, and question formats.
@@
We compare the ability of the MedMax model and the base model to caption as well as generate biomedical images for diverse domains. In total, we collect $1200$ instances from the testing split of PMC-OA (400), MIMIC-CXR (400), and unseen split of Quilt-1M (400) datasets. In particular, half of the dataset will be used for captioning evaluation and the other half will be used for generation evaluation. Similar to chambon2022roentgen [7], we extract the summary of the findings (impressions) from the report data and treat them as the ground-truth captions for the associated chest radiographs from MIMIC-CXR. This will highlight the medical report understanding and report-conditioned image generation capability of our model. Subsequently, we compute the BioMedCLIPScore to assess the closeness between the input image (caption) and predicted caption (image). We present the details for model inference in Appendix E.2.
Biomedical Visual Chatbot: We use the visual chatbot evaluation from LLaVA-Med consisting $193$ novel questions about $50$ unseen biomedical images. Specifically, the questions belong to two category: conversation and detailed description of the images. Subsequently, an LLM scores the predicted answer and the GPT-4 written reference answer out of 10 conditioned on the question, image caption and additional image context. Finally, we compute the average relative prediction score as the ratio of the score for predicted answer and score for the reference answer.
Biomedical Multimodal Generation: We utilize $500$ hidden instances of the MedMax-Instruct data for model evaluation. For a given text query, we prompt the base (or finetuned) model to generate interleaved response. Subsequently, we compare the text content in the predicted multimodal response with the reference text response using LLM (same as visual chat). In addition, we compare the generated image with the reference image using the image-image similarity score from the BioMedCLIP model. We present the evaluation templates and inference details in Appendix E.2. Further, we conduct contamination analysis and find that there are no exact matches between image-text pairs across these datasets. We provide the summary of the tasks and metrics in Table 1.
@@ 5 Experiments
Table 2: Performance of the MedMax model and baselines on the downstream VQA tasks. We find that the MedMax mixed-modal model outperforms closed as well as open multimodal models on twelve VQA datasets. This highlights that the model can generalize well to unseen instances and tasks ranging across biomedical domains.
@@ 5.1 Main results
Biomedical Visual Question Answering.
@@
We compare the performance of the finetuned MedMax model against several vision-language models on the battery of the VQA datasets in our evaluation suite. Specifically, the baselines includes the open models: the mixed-modal Chameleon-7B [34], and multimodal input but text-only biomedical specific models: LLaVA-Med-v1.5 [30], and HuatuoGPT-Vision-7B [64]. In addition, we consider two closed vision-language models: GPT-4o-mini and GPT-4o. We present our results in the Table 2. Our experiments reveal that MedMax model outperforms the base model (Chameleon) by $26.1\%$ percentage points on the average VQA performance across twelve tasks. This highlights that MedMax is suitable for building biomedical specialized model from the base model via instruction-tuning. Moreover, we observe that it achieves the best average performance across diverse baselines. In particular, it beats the closed vision model, GPT-4o, by $18.3\%$ percentage points. It indicates that the MedMax model is the most capable multimodal model for biomedical VQA across open as well as closed models. Further, we highlight the task-specific performance on individual VQA tasks too.
@@
In addition, we observe that MedMax achieves the highest accuracy amongst the baselines for 7 out of 12 tasks. Interestingly, we find that Chameleon-7B achieves the highest QuiltVQA (Closed) but a low QuiltVQA (Open) accuracy which suggests that closed-ended questions are easier to solve for the base model. Additionally, MedMax achieves $99.5\%$ performance on the unseen split of OmniMedVQA, demonstrating that this dataset can be solved with high accuracy when exposed to in-distribution examples. The model demonstrates the ability to answer questions about diverse anatomical regions after being exposed to some in-distribution data from OmniMedVQA. Furthermore, HuatuoGPT-Vision-7B outperforms other baselines on 3 out of 12 datasets. This can be attributed to its high-quality fine-tuning dataset and the Qwen-2 backbone [60], pre-trained on 7 trillion text tokens, which enhances the model’s biomedical knowledge and its ability to interpret diverse formatting styles.
@@
Additionally, we observe that the MedMax model achieves performance improvements of $10.7\%$, $15\%$, and $23\%$ on the unseen QuiltVQA (Open), PathMMU, and ProbMed datasets, respectively. This underscores the MedMax model’s ability to generalize effectively to novel downstream tasks through exposure to expert knowledge, question formats, and biomedical domains via instruction-tuning. Furthermore, we find that the MedMax is competitive to the task-specific finetuning of the LLaVA-Med on VQA-RAD, SLAKE and PathVQA datasets (Appendix F). This indicates that the practitioners can utilize an unified MedMax model instead of hosting separate task-specific LLaVA-Med models. Overall, these results highlight at the high-quality of the MedMax dataset for finetuning the mixed-modal foundation models for biomedical assistants.
Biomedical Image captioning and generation.
Here, we compare the capability of the MedMax model and base model to interpret and generate biomedical image data across different biomedical domains. We present the results in Figure 7. Our empirical findings suggest that the MedMax consistently outperforms the base model on the biomedical image captioning as well as the image generation task. In particular, the MedMax outperforms Chameleon by achieving a relative gain of $28\%$, $33\%$ and $14\%$ on biomedical captioning for the images in the PMC-OA, Quilt, and MIMIC-CXR datasets, respectively. Additionally, the MedMax also outperforms the Chameleon by achieving a relative gain of $100\%$, $14\%$, and $50\%$ on the image generation tasks across PMC-OA, Quilt,and MIMIC-CXR captions, respectively. Further, we compare the captioning abilities of our model against other models in Appendix H. Overall, our results demonstrate that the MedMax model excels at reasoning about images and generating realistic biomedical visuals aligned with user intent.
Biomedical Multimodal Generation.
(a) Evaluating text content.
(b) Evaluating image content.
Figure 6: Performance on the multimodal generation task. Comparison between the performance of the MedMax and Chameleon mixed-modal model on the multimodal generation task. We find that MedMax finetuning improves the multimodal content generation capabilities for the biomedical domain.
Here, we study the capability of the MedMax model to generate multimodal (interleaved image-text) content (Figure 6). Our empirical findings highlight that the MedMax outperforms Chameleon by achieving a relative improvement of $25.2\%$ on the quality of the text content in the multimodal output, as measured by LLM score. Furthermore, we observe that the MedMax outperforms Chameleon by achieving a relative improvement of $31.5\%$ on the synthesized biomedical image in the multimodal output, as measured by the image-image similarity score using BioMedCLIP. This indicates that instruction tuning with MedMax data equips a mixed-modal model with strong multimodal generation capabilities in the biomedical domain. Overall, our results lay the foundation for future exploration in unlocking and evaluating multimodal generation capabilities via instruction-tuning.
@@
Figure 8: Performance on the visual chat task. We find that the chatting capabilities of our model is quite competitive, suggesting its ability to answer novel queries about biomedical images.
Biomedical Visual Chat.
@@
We study the ability of the MedMax model and the baselines to answer novel queries about the biomedical images. We present the results in Figure 8. We observe that the overall LLM score of MedMax model is higher than the base Chameleon model and LLaVA-med-v1 by $34\%$ and $10.2\%$ percentage points. This indicates instruction-tuning enables strong visual chatting capability to the mixed-modal models. In addition, the MedMax model is quite competitive in comparison to the best-performing model, LLaVA-Med-v1.5, with a difference of $2.1\%$ percentage points. We observe that LLaVA-Med-v1.5 excels in the conversation split, while HuatuoGPT-Vision performs better on the description split of the dataset. These improvements are likely to stem from their high-quality language backbones, Mistral-v0.2-Instruct and Qwen-2, which enhance query interpretation and text generation. Future advancements in mixed-modal models will further improve biomedical visual chat performance with the MedMax.
5.2 Qualitative Examples
We present a few qualitative examples for MedMax-7B generations in Figure 12 and 13. The good qualitative examples highlight the model’s strong understanding of biomedical images and its ability to generate detailed answers across various tasks, especially in producing multimodal reports in response to questions. However, the examples also demonstrate the model’s limitations, including its occasional misinterpretation of image details and generation of images with meaningless textual artifacts.
6 Ablation Studies
@@ 6.1 Data scaling
6.2 Impact of Specific Data Subsets
@@
6.3 Impact of Specialized Visual Encoder
[…]
figurec
@@ 8 Conclusion
We present MedMax, the first instruction-tuning dataset enabling interleaved content generation for biomedical AI through MedMax-Instruct. MedMax also supports biomedical VQA, dialog, captioning, generation, and report understanding. Our experiments demonstrate that MedMax-tuned models achieve strong performance across multiple tasks, laying a strong foundation for the next generation of multimodal biomedical AI assistants. We emphasize that the MedMax model is a research prototype intended to foster community involvement in developing capable biomedical assistants. In medical emergencies, users should always consult a healthcare professional rather than relying solely on the model’s output. With continuous community feedback and the collection of high-quality expert data, we aim to enhance the model’s accuracy and safety over time.